# Collapse of wave functions in Schrödinger’s wave mechanics

### 25 Jan 2024 v1

@@ Abstract
We show that inelastic scattering can lead to a collapse of the wave function within standard evolution through the Schrödinger equation. The operational definition of collapse only requires that the initial width of the emerging wave function in inelastic scattering is primarily determined by the size of the participating scattering center, but not by the width of the incoming wave function. However, we also point out that collapse through inelastic scattering cannot replace the Born rule in quantum mechanics, but it leads to a more concise formulation of the measurement problem.
@@
Furthermore, the spatial extension of the signal from an individual particle will be determined by detector resolution, but not by the width of the probability density $|\psi(\bm{x},t)|^{2}$. This agrees with observations in single-particle diffraction experiments rf:merli ; rf:tonomura ; rf:bach . Interference effects generate macroscopic separations between different maxima of $|\psi(\bm{x},t)|^{2}$ and sampling the signals from many particles produces the interference patterns. However, every single particle signal is still pointlike in the sense of being determined by detector resolution.
@@
This led to the speculations about a special role of observers in the old Copenhagen interpretation of quantum mechancis: The combination of scattering and observation does not register any pre-existing orbit of the particle but creates that orbit rf:wh1 ; rf:bohr ; rf:bohr2 ; rf:wh2 ; rf:jvn . Heisenberg took a softer stance later in life and argued that localization of the particle during the measurement process may also be a consequence of complexity of interactions of the particle during the detection, see Heisenberg’s remarks in rf:wh3 , in particular p. 23.
@@
The assumption that localization happens due to complexity of interactions during the measurement process appears much more agreeable and contributed to the motivation for the decoherence program, which made important contributions to our understanding of the quantum-to-classical transition and the emergence of classically preferred states rf:zeh1 ; rf:zurek ; rf:zeh4 ; rf:phz ; rf:schlosshauer . However, complexity of interactions cannot solve the measurement problem as such. The interaction of an electron with the fluorescent screen in a low energy electron diffraction (LEED) experiment is no more complex than its coherent interaction with the surface atoms in the sample. Both kinds of interactions are described in order $e^{2}$ through first-order scattering amplitudes of an electron interacting with a medium, and there is no physical rationale for arguing that scattering off the fluorescent screen should be more complex than scattering off the sample. Indeed, the designation “measurement problem” is a misnomer that reflects the fact that particle-like behavior was perceived to be linked to observations in the early development of quantum mechanics. However, this is clearly not the case. Once we trigger the electron gun in an LEED device, the electrons will scatter coherently off the sample and produce pointlike flashes on the screen, even when we do not watch nor turn on a CCD camera for recording.
@@
Particle detection through inelastic scattering implies that we deliberately generate situations where inelastic scattering is likely to occur. However, the premise of science to identify laws of nature that desribe the universe independently from any conscious measurement or observation, implies that a particle will behave in exactly the same way irrespective of whether a particular medium is used for particle detection or not. I would posit that most scientists believe that a charged particle in a supersaturated vapor will generate a track in every corner of the universe, irrespective how far away from the next life form or civilization. The question is then not whether observation or measurement changes smooth evolution according to the Schrödinger equation, but whether the Schrödinger equation itself can explain the emergence of particle tracks.
Mott rf:mott and Heisenberg rf:wh2 had pointed out that forward domination of Coulomb scattering can naturally extend a particle track where a track has been started. However, as Schonfeld correctly noted, the key question from the point of view of the measurement problem is the formation of the beginning of the track, the first droplet rf:jfs2 . Based on properties of polarized molecules in supersaturated vapors, Schonfeld argues that Schrödinger evolution of the wave function should be able to explain the emergence of the track through a “consumption” or “draining” of the initial spherical wave through the scattered wave function.
However, I would submit that every medium that can generate repeated inelastic particle scattering, should generate a particle track, and this should not critically depend on properties of the scattering centers. We therefore model the medium as a set of scattering centers at locations $\bm{a}_{I}$, where $I$ enumerates the scattering centers. Furthermore, we assume fixed location of the scattering centers, as in a solid matrix. We are not aiming for realistic modeling of liquid-medium detectors. We are rather interested in a proof of principle investigation whether inelastic scattering can trigger wave function collapse within the scope of Schrödinger’s wave equation.
@@
The Hamiltonian would evolve an initial state $|\Psi(t^{\prime})\rangle=\exp(-\,\mathrm{i}H_{0}t^{\prime}/\hbar)|\Psi_{i}\rangle$ into the state $|\Psi(t)\rangle=\exp(-\,\mathrm{i}H_{0}t/\hbar)|\Psi_{f}\rangle$, where both initial and final states are expressed in terms of unperturbed states at initial time $t^{\prime}$ and final time $t$, respectively. Note that expressing the state at times $t^{\prime}$ and $t$ in terms of unperturbed states at time 00 per se does not require asymptotic conditions. It just defines the states $|\Psi_{i}\rangle$ and $|\Psi_{f}\rangle$ to which $|\Psi(t^{\prime})\rangle$ and $|\Psi(t)\rangle$ would be related in the absence of scattering.
@@
Free time evolution generates dispersion of wave packets with a time scale $\tau=2m\Delta x^{2}/\hbar$. However, the free dispersion time scale is much larger than particle travel times in lab experiments with electrons from millimetre aperture electron guns or neutrons from centimetre aperture neutron pipes, see, e.g. rf:rdqm3ed . Our condition for emergence of a pointlike signal from wave function collapse is therefore that the width of $\langle\bm{x}|\psi_{f}\rangle$ is primarily determined by the size of the scattering centers, but not by the width of $\langle\bm{x}|\psi_{i}\rangle$.
@@
Projection of Eq. (9) yields the outgoing wave packet $\psi_{f}(\bm{x})$ in terms of the incoming wave packet $\psi_{i}(\bm{x})$ (both taken at fiducial time $t=0$),
@@
The Dirichlet kernel $\sin\!\left(\kappa|\bm{z}-\bm{x}^{\prime}|\right)/(\pi|\bm{z}-\bm{x}^{\prime}|)$ in Eq. (11) provides an Ångström-scale approximation to a $\delta$ function already for $k=0$ and for particle masses $m\geq m_{e}=511\,\mathrm{keV}$ if the excitation energy satisfies $\hbar\omega_{10}\gtrsim 1\,\mathrm{eV}$. Furthermore, the factor $\phi^{+}_{1}(\bm{y}-\bm{a}_{I})\phi_{0}(\bm{y}-\bm{a}_{I})$ will provide atom-scale resolution around the location $\bm{a}_{I}$ of the scattering center, and screened Coulomb potentials have Ångström-scale range in dense materials. All this indicates that the emerging wave packet from inelastic scattering should be located at the scattering center and have an initial width that is determined by the width of the scattering center, in perfect agreement with the operational definition of a particle-like signal. Indeed, approximating the Ångström-scale factors in Eq. (11) with $\delta$-functions yields
@@
However, this does not absolve us of the measurement problem in wave mechanics. The catch is that nothing in the mathematical evolution of the wave function through inelastic scattering prevents the superposition of two inelastic scattering events in locations $\bm{a}_{1}$ and $\bm{a}_{2}$. This will yield the superposition of a pointlike particle signal in location $\bm{a}_{1}$ due to excitation of the scattering center $\bm{a}_{1}$, with another pointlike particle signal in location $\bm{a}_{2}$ due to excitation of the scattering center $\bm{a}_{2}$. This is an allowed final state of the scattering matrix, just like the individual pointlike signal states are allowed final states. Within the purely mathematical evolution of the wave function according to the Schrödinger equation, collapse of part of the extended wave function of the incoming particle around one atom on a fluorescent screen or one molecule in a supersaturated vapor does not preclude simultaneous collapse of another part of the wave function around another atom or molecule. With respect to the detector, this generates a superposition of states proportional to $|\phi_{1,\bm{a}_{1}}\rangle|\phi_{0,\bm{a}_{2}}\rangle+\exp(\mathrm{i}\beta)|\phi_{0,\bm{a}_{1}}\rangle|\phi_{1,\bm{a}_{2}}\rangle$ (up to normalization, and with an arbitrary relative phase $\beta$). Without the Born rule, the Schrödinger equation inevitably allows for the formation of more than one particle track, or more than one point lit on a fluorescent screen.
Obviously, the Born rule amounts to a superselection rule stating that the incoming particle can excite only one atom or molecule in a medium at a time, and the pointlike particle signal is then a consequence of the behavior (12) in inelastic scattering. Notwithstanding, the measurement problem of Schrödinger’s wave mechanics persists in the form: The Schrödinger equation on its own cannot explain the breakdown of the superposition principle in inelastic particle scattering in a medium.
We acknowledge that epistemic interpretations of the wave function rf:rovelli1 ; rf:epi1 ; rf:epi2 ; rf:rovelli2 ; rf:epi3 or Bohmian mechanics rf:bohm ; rf:duerr ; rf:maudlin ; rf:goldstein do not suffer from this problem. However, completely surrendering a realist interpretation of the wave function, or modifying the Schrödinger equation (which is the universal nonrelativistic limit of relativistic wave equations), are high prizes to pay.

### 01 Mar 2024 v2

@@ Abstract
We show that inelastic scattering leads to a collapse of the wave function within standard evolution through the Schrödinger equation, whereas elastic scattering will not collapse the wave function. Specifically, we find that the initial width of the emerging wave function in inelastic scattering is primarily determined by the size of the participating scattering center, but not by the width of the incoming wave function. This implies that dynamical collapse of the wave function through inelastic scattering, together with energy quantization in bound quantum systems, can explain the emergence of particle-like signals without the need to invoke the Born rule.
@@
Furthermore, the spatial extension of the signal from an individual particle will be determined by detector resolution, but not by the width of the probability density $|\psi(\bm{x},t)|^{2}$. This agrees with observations in single-particle diffraction experiments [1, 2, 3]. Interference effects generate macroscopic separations between different maxima of $|\psi(\bm{x},t)|^{2}$ and sampling the signals from many particles produces the interference patterns. However, every single particle signal is still pointlike in the sense of being determined by detector resolution.
@@
This led to the speculations about a special role of observers in the old Copenhagen interpretation of quantum mechancis: The combination of scattering and observation does not register any pre-existing orbit of the particle but creates that orbit [4, 5, 6, 7, 8]. Heisenberg took a softer stance later in life and argued that localization of the particle during the measurement process may also be a consequence of complexity of interactions of the particle during the detection, see Heisenberg’s remarks in [9], in particular p. 23.
@@
The assumption that localization happens due to complexity of interactions during the measurement process appears much more agreeable and contributed to the motivation for the decoherence program, which made important contributions to our understanding of the quantum-to-classical transition and the emergence of classically preferred states [10, 11, 12, 13, 14]. However, complexity of interactions cannot solve the measurement problem as such. The interaction of an electron with the fluorescent screen in a low energy electron diffraction (LEED) experiment is no more complex than its coherent interaction with the surface atoms in the sample. Both kinds of interactions are described in order $e^{2}$ through first-order scattering amplitudes of an electron interacting with a medium, and there is no physical rationale for arguing that scattering off the fluorescent screen should be more complex than scattering off the sample. Indeed, the designation “measurement problem” is a misnomer that reflects the fact that particle-like behavior was perceived to be linked to observations in the early development of quantum mechanics. However, this is clearly not the case. Once we trigger the electron gun in an LEED device, the electrons will scatter coherently off the sample and produce pointlike flashes on the screen, even when we do not watch nor turn on a CCD camera for recording.
@@
Particle detection through inelastic scattering implies that we deliberately generate situations where inelastic scattering is likely to occur. However, the premise of science to identify laws of nature that desribe the universe independently from any conscious measurement or observation, implies that a particle will behave in exactly the same way irrespective of whether a particular medium is used for particle detection or not. We would posit that most scientists believe that a charged particle in a supersaturated vapor will generate a track in every corner of the universe, irrespective of how far away from the next life form or civilization. The question is then not whether observation or measurement changes smooth evolution according to the Schrödinger equation, but whether the Schrödinger equation itself can explain the emergence of particle tracks.
Mott [15] and Heisenberg [7] had pointed out that forward domination of Coulomb scattering can naturally extend a particle track where a track has been started. However, as Schonfeld correctly noted, the key question from the point of view of the measurement problem is the formation of the beginning of the track, the first droplet [16]. Based on properties of polarized molecules in supersaturated vapors, Schonfeld argues that Schrödinger evolution of the wave function should be able to explain the emergence of the track through a “consumption” or “draining” of the initial spherical wave through the scattered wave function.
However, we would submit that every medium that can generate repeated inelastic particle scattering, should generate a particle track, and this should not critically depend on properties of the scattering centers. We therefore model the medium as a set of scattering centers at locations $\bm{a}_{I}$, where $I$ enumerates the scattering centers. Furthermore, we assume fixed location of the scattering centers, as in a solid matrix. We are not aiming for realistic modeling of liquid-medium detectors. We are rather interested in a proof-of-principle investigation whether inelastic scattering can trigger wave function collapse within the scope of Schrödinger’s wave equation.
@@
The Hamiltonian would evolve an initial state $|\Psi(t^{\prime})\rangle=\exp(-\,\mathrm{i}H_{0}t^{\prime}/\hbar)|\Psi_{i}\rangle$ into the state $|\Psi(t)\rangle=\exp(-\,\mathrm{i}H_{0}t/\hbar)|\Psi_{f}\rangle$. Here both the initial and the final state of the particle-plus-medium system are expressed in terms of fiducial states $|\Psi_{i}\rangle$ and $|\Psi_{f}\rangle$ at time 00, where the fiducial states are related to the actual system state at times $t^{\prime}$ and $t$ through free time evolution, i.e. we describe the state before and after inelastic scattering through asymptotic free states. This complies with the assumption of finite range of the scattering potential $V$.
@@
Free time evolution generates dispersion of wave packets with a time scale $\tau=2m\Delta x^{2}/\hbar$. However, the free dispersion time scale is much larger than particle travel times in lab experiments with electrons from millimetre aperture electron guns or neutrons from centimetre aperture neutron pipes, see, e.g. [17]. Our condition for emergence of a pointlike signal from wave function collapse is therefore that the width of $\langle\bm{x}|\psi_{f}\rangle$ is primarily determined by the size of the scattering centers, but not by the width of $\langle\bm{x}|\psi_{i}\rangle$.
@@
Evaluation of Eq. (9) in the usual limits $t\to\infty$, $t^{\prime}\to-\,\infty$ (which practically only means that $t-t^{\prime}$ should be large compared to the travel time of the particle through the range of the scattering potential), yields the outgoing wave packet $\psi_{f}(\bm{x})$ in terms of the incoming wave packet $\psi_{i}(\bm{x})$, both taken at fiducial time 00,
@@
The Dirichlet kernel $\sin\!\left(\kappa|\bm{z}-\bm{x}^{\prime}|\right)/(\pi|\bm{z}-\bm{x}^{\prime}|)$ in Eq. (11) provides an Ångström-scale approximation to a $\delta$-function already for $k=0$ and for particle masses $m\geq m_{e}=511\,\mathrm{keV}/c^{2}$ if the excitation energy satisfies $\hbar\omega_{10}\gtrsim 1\,\mathrm{eV}$. Furthermore, the factor $\phi^{+}_{1}(\bm{y}-\bm{a}_{I})\phi_{0}(\bm{y}-\bm{a}_{I})$ will provide atom-scale resolution around the location $\bm{a}_{I}$ of the scattering center, and screened Coulomb potentials have Ångström-scale range in dense materials. All this indicates that the emerging wave packet from inelastic scattering should be located at the scattering center and have an initial width that is determined by the width of the scattering center, in perfect agreement with the operational definition of a particle-like signal. Indeed, approximating the Ångström-scale factors in Eq. (11) with $\delta$-functions yields
@@
We note that this derivation of wave function collapse works only for inelastic scattering. The scattered wave function for elastic scattering will always be a superposition of the 0th order term $\psi^{(0)}_{f}(\bm{x})\sim\psi_{i}(\bm{x})$ and a first order term $\psi^{(1)}_{f}(\bm{x})$. The 0th order term implies that the elastically scattered wave packet should at least inherit the width of the incoming wave packet. Furthermore, the Dirichlet kernel for elastic scattering contains no energy transfer term $2m\Delta E/\hbar^{2}$ in the argument, and this implies that the Dirichlet kernel will not provide an atom-scale approximation to a $\delta$-function for small particle momenta. The scattering matrix tells us that wave function collapse will happen in inelastic scattering, but not in elastic scattering.
However, wave function collapse in inelastic scattering by itself may not absolve us of the measurement problem in wave mechanics. The catch is that the mathematical evolution of the wave function through inelastic scattering might also allow for the superposition of two inelastic scattering events in locations $\bm{a}_{1}$ and $\bm{a}_{2}$. This would yield the superposition of a pointlike particle signal in location $\bm{a}_{1}$ due to excitation of the scattering center $\bm{a}_{1}$, with another pointlike particle signal in location $\bm{a}_{2}$ due to excitation of the scattering center $\bm{a}_{2}$. In the absence of any additional restrictions on permissible final states, coherent superposition of two inelastic scattering events appears to correspond to an allowed final state of the scattering matrix. With respect to the detector, this would generate a superposition of states proportional to $|\phi_{1,\bm{a}_{1}}\rangle|\phi_{0,\bm{a}_{2}}\rangle+\exp(\mathrm{i}\beta)|\phi_{0,\bm{a}_{1}}\rangle|\phi_{1,\bm{a}_{2}}\rangle$ (up to normalization, and with an arbitrary relative phase $\beta$). Without the Born rule, the Schrödinger equation might allow for the formation of more than one particle track, or more than one point lit on a fluorescent screen.
This is not a problem in epistemic interpretations of the wave function [18, 19, 20, 21, 22, 23] because the Born rule is understood as a necessary prescription for translating an epistemic wave function into a position signal.
It is also not a problem in Bohmian mechanics [24, 25, 26, 27] because the inclusion of a quantum potential and the addition of a guiding equation make the Born rule redundant.
Notwithstanding, wave function collapse in inelastic scattering opens up a third line of attack on the measurement problem. We may be able to exclude the simultaneous excitation of different scattering centers through the observation that low-energy excitations in isolated bound quantum systems can only yield energy eigenstates. This excludes states of the form [^1] We assume that we can neglect interactions between different scattering centers. $|\phi_{1,\bm{a}_{1}}\rangle|\phi_{0,\bm{a}_{2}}\rangle+\exp(\mathrm{i}\beta)|\phi_{0,\bm{a}_{1}}\rangle|\phi_{1,\bm{a}_{2}}\rangle$ as final detector states and explains the emergence of a single pointlike position signal without invoking the Born rule.